\begin{thebibliography}{999}
\bibitem{ALILI et al.}Alili L., Matsumoto H. and Shiraishi T.,  On a triplet of exponential
Brownian functionals,  S\'eminaire de Probabilit\'es XXXV, 396–415,
Lecture Notes in Math., 1755, Springer-Verlag. MR1837300 (2001).
\bibitem{ABDELHAY et al.} Abdelhaye Y. O. M.,  Badahi O. M.
and Ould Moustapha M. V., Wave kernel for the Schr\"odinger operator with the Morse potential and applications,
Far East Journal of Mathematical Sciences (FJMS) 102, Number 7, 1523-1532(2017).
%\bibitem{AYAZ-INTISSAR} Ayaz K., Intissar A., Selberg trace formulae for heat and wave kernels of Maass Laplacians on compact forms of the
%complex hyperbolic space $H^n(\C)$, no 2. Differential Geom. Appl. 15 , no. 1, 1–31 (2001).
%\bibitem{BOUSSEJRA-INTISSAR} Boussejra, A. and Intissar, A, $L^2-$concrete spectral analysis of the invariant 
%Laplacians $\Delta_{\alpha\beta}$ in the unit complex ball $B_n$, Journal of functional analysis 160, 115-140 (1998).
%\bibitem{BOUNK ET al.} Bunke, U. Olbrich, M. and Juhl, A. The wave kernel for the Laplacian on locally 
%symmetric spaces of rank one, Theta functions, Trace formulas and the Selberg zeta 
%function, Ann. Global Anal. Geom. {\bf 12}  357-405(1994).
\bibitem{COMTET} Comtet A., On the Landau levels on the hyperbolic plane, Ann. Physics 173, 185–209(1987).
%\bibitem{2} \label{2} V. Ditkine et A. Proudnikov, Transformations
% integrales et calcul op\'erationel, Traduction francaise
% Edition MIR 1978.
\bibitem{ELSTRODT} Elstrodt J., Die Resolvente zum Eigenwertproblem da automorphen formen in der 
hyperbolishen Ebene I, Math. Ann. 203,  295-330(1973).
\bibitem{ERDELYI et al.} Erdelyi  A,  Magnus W., Oberhettinger F. and Tricomi F., Higher transandental function Vol. I, McGraw-Hill New York 1953.
%\bibitem{EXTON} Exton H., Multiple Hypergeometric functions and applications
%Chichestyer, New York-London-Sydney-Toronto (1976).
\bibitem{FAY}  Fay J., Fourier coefficient of the resolvents for fuschien groups Reine Engew math. 293,143-203(1977).
%\bibitem{FERAPONTOV-VESEL} E. V. Ferapontov and A. P. Vesel,   Integrable Schrödinger operators with magnetic fields: Factorization method on curved
%surfaces, 
%Journal of Mathematical Physics 42, 590 (2001) doi: 10.1063/1.1334903.
%\bibitem{FIDIANI} Fidiani E., Modeling of Diatomic molecule using the Morse Potential and the Verlet algorithm, AIP conferences Procedings1719030001(2016)
%doi 10-1063/1.4943696.
%\bibitem{GHANMI-INTISSAR} Ghanmi A., Intissar A., Asymptotic of complex hyperbolic geometry and L2-spectral analysis of Landau-like Hamiltonians.
%J. Math. Phys. 46, no. 3, 032107(2005).
%\bibitem{HASSANABADI-ZARE} Hassanabadi H.,  Zare S.,
Investigation of quasi-Morse potential in position-dependent
mass formalism,
Eur. Phys. J. Plus  132: 49  DOI 10.1140/epjp/i2017-11319-x(2017).
\bibitem{IKEDA-MATSUMOTO}  Ikeda N., Matsumoto H.,  Brownian motion
one the hyperbolic plane and Selberg trace formula, J. Funct. Anal. 163, 63-110(1999).
% \bibitem{INTISSAR-MOUSTAPHA-1} Intissar A.,  Ould Moustapha. M. V.,  Solution explicite de l'\'equation des ondes dans un espace sym\'etrique de type non compact de rang 1, C.R.Acad. Sci. Paris 321, 77-81(1995).
% \bibitem{INTISSAR-MOUSTAPHA-2}   Intissar A. and
% Ould Moustapha M. V., Explicit formulas for the Wave Kernels for the Laplacian $\Delta_{\alpha\beta}$ on the Bergman ball $B^n$, $n\geq 1$, Annals of Global Anal. and Geometry 15: 221-234, (1997).
\bibitem{LEBEDEV}  Lebedev, N., N., Special Functions
and their applications; Dover Publications INC New York (1972).
\bibitem{MAASS}  Maass M., Die differentialglichungen in der theorie der elliptischen modufunktionen,
Math. Ann. 125, 235-263(1953).
\bibitem{MAGNUS et al.} Magnus, F. Oberhettinger and R.P. Soni, Formulas and Theorems for special functions of Mathematical Physics, Third enlarged edition
,Springer-Verlag Berlin-Heidelberg-New York (1966).
\bibitem{MORSE} Morse P. M., Diatomic molecules according to the wave mechanics. II. Vibrational levels,
Phys. Rev. 34, 57-64(1929).
% \bibitem{OULD MOUSTAPHA1} Ould Moustapha M. V., Slutions explicites des \'equations des ondes et de la chaleur pour les Laplaciens $\Delta_{\alpha\beta}$ dans la boule unit\'e de $C^n$, Th\`ese Doctorat d'Etat, Universit\'e Mohamed V, Facult\'e
% des Sciences de Rabat (1997).
\bibitem{OULD MOUSTAPHA2} Ould Moustapha  M. V.,  Wave kernels with magnetic field on the hyperbolic plane and
with the Morse potential on the real line,
Quantum Stud.: Math. Found.
https://doi.org/10.1007/s40509-019-00200-x(2019).
\bibitem{OULD MOUSTAPHA3} Ould Moustapha M. V., Resolvent kernels with magnetic fields on the hyperbolic plane and applications, https://arxiv.org/abs/1909.13487.
\bibitem{PRUDNIKOV et al.} A. P. Prudnikov, Y. A. Brichkov, O. I. Marichev, Integrals and series, Vol. 5, Inverse Laplace transforms,
Gordon and Breach Sciences publishers, New York-Paris Tokyo-Melborne-Montreux,(1992).
\bibitem{ROELCKE} Roelcke W. Das Eigenwertproblem der automorphen Formen in der hyperbolischen
Ebene, I, Math. Ann. 167 (1966), 292-337; II, Math. Ann. 168, 261-324(1967).
%\bibitem{TASSELI}  Tasseli H., Exact solutions for vibrational
%Levels of Morse potential, J.Phys. A: Math.Gen.
%31, 779-788(1998).
%\bibitem{ZNOJIL} Znojil  M.,  Morse Potential, Symmetric Morse Potential and Poracket bound-states energies, Mod. Phys. Letter A31 (2016).
 \end{thebibliography}

\documentclass[12pt]{article}
\usepackage{amsmath,amsfonts,amssymb,amsthm, epsfig,
epstopdf, titling, url, array}
\usepackage[utf8]{inputenc}
\usepackage{mathtools}
\usepackage[english]{babel}
\usepackage{hyperref}
\newtheorem{thm}{Theorem}[section]

\newtheorem{cor}{Corollary}[section]
\newtheorem{prop}{Proposition}[section]
\numberwithin{equation}{section}

\def\C{\mathbb C}
\def\D{\mathbb D}

\def\R{I\!\!R}
\def\H{I\!\!H}
\def\C{I\!\!\!\!C}

\def\D{I\!\!D}
\title{Resolvent kernels and time dependent Schr\" odinger equations under magnetic field on the hyperbolic half plane $\H$ and the Morse potential on the real line $\R$ }
\author{Mohamed Vall Ould Moustapha}

\begin{document}
\maketitle
\begin{abstract}

This paper deals with exact  formulas for the resolvent kernels and exact solutions of time dependent  Schr\"odinger equations under a uniform magnetic field on the hyperbolic half plane $\H$, and under a diatomic molecular Morse potential on the real line $\R$. 
\end{abstract}
 
\section{Introduction}
 
In this paper we give an explicit formulas for the Schwartz integral kernels $\widetilde{G}_k(\lambda, z,  z')$, ${\cal G}^{\lambda}_k(\mu, X,  X')$ , $\widetilde{H}_k(\lambda, z,  z')$ and ${\cal H}^{\lambda}_k(\mu, X,  X')$ of the resolvent and evolution operators 
 $\left(\widetilde{{\cal D}}_k+\lambda^2\right)^{-1}$, $\left(\Lambda^{\lambda, k}+\mu^2\right)^{-1}$, $e^{t \widetilde{{\cal D}}_k}$  and $e^{t \Lambda^{\lambda, k}}$.

The operators $\widetilde{{\cal D}}_k$ and $\Lambda^{\lambda, k}$ are
the modified Schr\"odinger operators with magnetic field  on the hyperbolic upper half plane $\H$  and the Schr\"odinger operator with the diatomic molecular Morse potential on the real line $\R$, they are respectively given by (Ikeda-Matsumoto \cite{IKEDA-MATSUMOTO})
\begin{align}\widetilde{{\cal D}}_{k}=y^2\left(\frac{\partial^2}{\partial x^2}+\frac{\partial^2}{\partial y^2}\right)+ 2 i k y\frac{\partial}{\partial x}+ \frac{1}{4},\end{align}
and
\begin{align}\label{Morse}\Lambda^{\lambda, k}=\frac{\partial^2}{\partial X^2}-2k\lambda e^X -\lambda^2 e^{2X}.\end{align}
Note that the operator $\widetilde{{\cal D}}_{k}$ is introduced firstly by Maass \cite{MAASS}  and can be written as
\begin{align}\widetilde{{\cal D}}_k= {\cal L}_{k}^{\H}+k^2+\frac{1}{4},\end{align} 
where ${\cal L}_{k}^{\H}$ is the  Schr\"odinger operator with uniform magnetic field on the hyperbolic upper half plane $\H$ 
\begin{align}{\cal L}_{k}^{\H}=-(z-\overline{z})^2)\frac{\partial^2}{\partial z\partial\overline{z}}+k(z-\overline{z})\left(\frac{\partial}{\partial z}+\frac{\partial
}{\partial \overline{z}}\right)-k^2.\end{align}


The operators  ${\cal L}_{k}^{\H}$   has a physical interpretation as being the Hamiltonian which governs a non relativistic charged particle
moving under the influence of the magnetic field of constant strength $|k|$, perpendicular to $\H$. The operator $\widetilde{{\cal D}}_{k}$  is basic to the work of Roelcke \cite{ROELCKE}, Elstrodt \cite{ELSTRODT}, Fay  \cite{FAY} and Comtet \cite{COMTET}.\\
 The purely vibrational levels of diatomic
molecules with angular momentum $l=0$ have been described by the Morse potential since $1929$ (see Morse\cite{MORSE}).\\
The operators  $\widetilde{{\cal D}}_{k}$ and $\Lambda^{\lambda, k}$ are,
non-positive, definite
and each of them has an absolute continuous spectrum and a points spectrum if $|k|\ge 1/2$ .\\
For a recent work on the magnetic field on the hyperbolic plane and the Morse Potential (see Ould Moustapha \cite{OULD MOUSTAPHA2} and the references therein).\\
 For $k=0$, the operator  $\widetilde{{\cal D}}_{0}$  reduces to the free Laplace-Beltrami operators on the hyperbolic half-plane and
the free resolvent kernel on the hyperbolic upper half plane 
is given by 
\begin{align*} \widetilde{G}_0(\mu, z, z')=
\frac{\Gamma(s)\Gamma(s)}{4\pi\Gamma(2s)}\cosh^{-2s}\rho( z,  z'){}_2F_1\left(s, s , 2s,  \cosh^{-2}\rho(z, z')\right),\end{align*}
where $s=(1-i\mu)/2$ and $\rho(z,  z')$ is  the hyperbolic half plane distance.\\
The function ${}_2F_1(a, b, c, z)$ is the Gauss hypergeometric function defined by:
\begin{align}\label{Gauss}
  {}_2F_1(a, b, c, z)=\sum_{n=0}^{\infty}\frac{(a)_n(b)_n}{(c)_n n!}z^n,
  \quad |z|<1,
\end{align}
  $(a)_n=\frac{\Gamma(a+n)}{\Gamma(a}$  is the Pochhamer symbol,
and $\Gamma$ is the classical Euler function.\\
For $k=0$ the operator $\Lambda^{\lambda, 0}$ reduces to the Schr\"odinger operator with the so-called Liouville potential. In this case the resolvent kernel is given by the explicit formula Ikeda-Matsumoto\cite{IKEDA-MATSUMOTO}
\begin{align}G^{M}_{\lambda, 0}(\mu, X, X')=2 I _{i\mu}(\lambda e^X)K_{i\mu}(\lambda e^{X'}).\end{align}
For the wave kernel of the operator $\Lambda^{\lambda, 0}$ see Abdelhaye et al. \cite{ABDELHAY et al.}\\
The organization of the paper is as follows. In sections 2 and 3 we give explicit formula for the resolvent kernel and we solve the time dependent  Schr\"odinger equation  with magnetic field 
on the hyperbolic half plane $\H$. Sections 4  and 5  are devoted to explicit formulas for the resolvent kernels and the solution of the time dependent  Schr\"odinger equation with diatomic molecular Morse
 potential on $\R$.
Finally in section 6 we introduce and discuss some applications.
\section{Resolvent kernel under the magnetic field on the hyperbolic half  plane $\H$ }
Let  $\H=\{z=x+iy\in \C,\, y>0\}$ be the upper half plane endowed with the usual hyperbolic metric $\widetilde{ds}=\frac{dx^2+dy^2}{y^2},$ 
the Riemannian space $(\H, \widetilde{ds})$ is called the hyperbolic upper half plane.
 The metric $\widetilde{ds}$ is invariant with respect to the group $\widetilde{G}=SL_2(\R)$ with
 $$ SL_2(\R)=\left\{ \left(
\begin{array}{cc}
a & b\\
c & d
\end{array}\right) a, b, c, d \in \R: a d-c b=1\right\}.$$
The hyperbolic distance between two points $z, z' $ is given by
\begin{align}\label{dist2} \cosh^2 (\rho(z, z')/2)=\frac{(x-x')^{2}+(y+y')^{2}}{4yy'}.\end{align} 
 It is well known that the hyperbolic upper half plane $(\H, \widetilde{ds})$ is isomorphic to the hyperbolic disc $(\D, ds)$ 
via the Cayley transform:
$$ z=c^{-1}w=-i\frac{w+1}{w-1},
w=c z=\frac{z-i}{z+i}. $$
Let
${\cal D}_{k}$ be  the modified Schr\"odinger operator with uniform magnetic field on the hyperbolic disc $\D$, defined by  (Ould Moustapha \cite{OULD MOUSTAPHA2})
 \begin{align*}{\cal D}_{k}=(1-|w|^2)^2\frac{\partial^2}{\partial w \partial \overline{w}}+k(1-|w|^2)w\frac{\partial}{\partial w} \nonumber\\
-k(1-|w|^2)\overline{w}\frac{\partial}{\partial \overline{w}}-k^2|w|^2+k^2+\frac{1}{4}. \end{align*}
The operators ${\cal D}_{k}$ and $\widetilde{{\cal D}}_{k}$ are related by the formula
\begin{align}\label{intertwin}U_{k}\widetilde{{\cal D}}_{k}U^{-1}_{k}f(c^{-1}w)={\cal D}_{k} f(c^{-1}w),\end{align}
with
\begin{align}\label{U}\left(U_{k}f\right)(w)=\left(\frac{1-\overline{w}}{1-w}\right)^{k}f(c^{-1}w),\ \ \left(U^{-1}_{k}g\right)(z)=\left(\frac{i-\overline{z}}{z+i}\right)^{k}g(c z).\end{align} 
\begin{prop}\label{Resolvent} The resolvent kernel associated to the modified Schr\"odinger operators with magnetic field $\widetilde{{\cal D}}_{k}$ on the hyperbolic half plane $\H,$
 is given by
$\widetilde{G}_k(s, z, z')=\frac{\Gamma(s-k)\Gamma(s+k)}{4\pi\Gamma(2s)}\left(\frac{z'-\overline{ z}}{z-\overline{z'}}\right)^{k}\left(\cosh^2\rho(z, z')/2\right)^{-s}\times $  \begin{align}F\left(s-|k|, s+|k|, 2s, \cosh^{-2}\rho(z, z')/2\right) 
\end{align} 
with $\cosh^2\rho(z, z')/2$ is given in \eqref{dist2}
 \end{prop}
 \begin{proof} Using the formula \eqref{U} 
we have 
\begin{align} \label{transport}
U_k^z [U_{-k}^{z'}[\widetilde{G}_k(t, z, z')](w')](w')=G_k(s, w, w')
\end{align}
where $G_k(s, w, w')$ is resolvent kernel with modified uniform magnetic potential on the hyperbolic disc given by (Ould Moustapha \cite{OULD MOUSTAPHA2})\\
$\label{G}G_k(s, w, w')=
\frac{\Gamma(s-k)\Gamma(s+k)}{4\pi\Gamma(2s)}\left(\frac{1-w\overline{w'}}{1-\overline{w}w'}\right)^{k}\times $ 
\begin{align}\left(\cosh^2 (d(w, w')/2\right)^{-s}F\left(s-|k|, s+|k|, 2s, \cosh^{-2}(d(w, w')/2\right),
\end{align}
 \begin{align}\label{dist1}\cosh^2 (d(w, w')/2)=\frac{|1-w\overline{w'}|^2}{(1-|w|^2)(1-|w'|^2)}, \end{align}
and from the formula \eqref{intertwin} 
we obtain the results of the proposition.
\end{proof}
 \begin{prop}  Let $\widetilde{G}_k(\mu, z, z')$ be the resolvent kernel with uniform magnetic field on the hyperbolic upper half plane, then  we have \\ 
\begin{align}\label{key}\widetilde{G}_k(\mu, z, z')=\int_\rho^{+\infty}\widetilde{W}_k(b,\rho)\frac{e^{-i\mu b}}{2i\mu} d b,  \end{align}
where\\
$\widetilde{W}_k(b, \rho)=\frac{1}{2\pi}\left(\frac{z'-\overline{z}}{z-\overline{z'}}\right)^{k}
(\cosh^2b/2-\cosh^2\rho/2)^{-1/2}\times $ \begin{align}\label{Kernel} F\left(|k|,-|k|, 1/2,1-\frac{\cosh^2b/2}{\cosh^2\rho/2}\right),\end{align}
and  $\rho(z, z')$  is the distance between two points $z$ and $z'$ given in  \eqref{dist1}.
\end{prop}
\begin{proof}
The proof of this proposition is a consequence of the formula \eqref{transport}  and the  Theorem 2.3 (Ould Moustapha \cite{OULD MOUSTAPHA3}).
\end{proof}

\begin{prop} Let $\widetilde{W}_k(b, \rho)$ be the kernel in \eqref{Kernel}, the following formulas hold\\
{\bf i)}
$\widetilde{W}_k(b, \rho)=\frac{1}{2\pi}\left(\frac{z'-\bar z}{z-\bar{z'}}\right)^{k}(\cosh^2(b/2) - 
\cosh^2(\rho(z, z')/2))_+^{-1/2}\times$\\
$F(2|k|,-2|k|;1/2,(1-\frac{\cosh (b/2)}{\cosh (\rho/2)}).$\\
{\bf ii)}
$\widetilde{W}_k(b, \rho)=\frac{1}{2\pi}\left(\frac{z'-\bar z}{z-\bar{z'}}\right)^{k}\left(\frac{\cosh^2b/2}{\cosh^2\rho/2}\right)^{|k|}(\cosh^2(b/2) - 
\cosh^2(\rho(z,z')/2))_+^{-1/2}\times$\\
$ F(-|k|,1/2-|k|;1/2,1-\frac{\cosh^2 (\rho/2)}{\cosh^2 (b/2)}).$\\
{\bf iii)} For $|k|$ integer or half integer we have:\\
$\widetilde{W}_k(b, \rho)=\frac{1}{2\pi}\left(\frac{z'-\bar z}{z-\bar{z'}}\right)^{k}(\cosh^2(b/2) - 
\cosh^2(\rho(z,z')/2))_+^{-1/2}T_{2|k|}\left(\frac{\cosh (b/2)}{\cosh (\rho(z,z')/2)}\right)$\\
where $T_{2|k|}(x)$ is the Chebichev polynomial of the first kind.\\
{\bf iv)}  For $|k|$ integer or half integer we have\\
$\widetilde{W}_k(b, \rho)=\frac{1}{2\pi}\left(\frac{z'-\bar z}{z-\bar{z'}}\right)^{k}\left(\cosh(\rho(z, z')/2)\right)^{-2|k|}\times$ \\  $\sum_{n=0}^{n=[|k|]}\frac{(-|k|)_n (1/2-|k|)_n}{(1/2)_n n!}\cosh(b/2)^{2|k|-2n}\left(\cosh^2b/2-\cosh^2\rho/2\right)^{n-1/2}.$\\
where $[|k|]$ is the entire part of $|k|$.
\end{prop}
\begin{proof}
Using the formula Magnus et al. \cite{MAGNUS et al.},$ p.50$\\
$F(a, b, (a+b+1)/2, z)=F(a/2, b/2, (a+b+1)/2, 4z-4z^2)$
we get i).\\ 
To see ii) we apply the following formula Magnus et al. \cite{MAGNUS et al.},$p. 47$\\
$F(a, b, c, z)=(1-z)^{-a}F(a, c-b, c, \frac{z}{z-1})$.\\
The part iii) is a consequence of the formula  Magnus et al. \cite{MAGNUS et al.}, $p.39$,\\
$T_n(1-2x)=F(-n, n, 1/2, x)$.\\
Finally iv) can be shown from ii) and  the formula  Magnus et al. \cite{MAGNUS et al.} p.$39$ 
$F(-m, b, c, z)=\sum_{n=0}^{n=[m]}\frac{(-m)_n (b)_n}{(c)_n n!}z^n$
and the Proof of Lemma is finished.\\
\end{proof}
\section{The time dependent Schr\"odinger equation under a magnetic field on the hyperbolic upper half plane}
Now, we give the solution of the following time dependent Schr\"odinger equation under a uniform magnetic field on the hyperbolic upper half plane. 
\begin{align}\label{Cauchy-Heat} \left \{\begin{array}{cc}\widetilde{\cal D}_k u(t, z)=\frac{\partial}{\partial
t}u(t, z)
, (t, z)\in \R^\ast_+\times \H \\ u(0, z)=0, u_t(0, z)=u_1(z), u_1\in
C^\infty_0(\H)\end{array}
\right., \end{align}
\begin{cor} The Schwartz integral  kernel of the heat operator $e^{t \widetilde{\cal D}_k}$  that solves the heat Cauchy problem\eqref{Cauchy-Heat}\ with uniform magnetic field on the hyperbolic plane is given by:
\begin{align}\label{heat kernel}H_k(t, z, z')=\int_r^\infty\frac{ e^{-b^2/4t}}{(4\pi t)^{3/2}}\widetilde{W}_k(b, z,  z')b db,\end{align}
where\\
$\widetilde{W}_k(r, \rho)=\frac{1}{2\pi}\left(\frac{z'-\overline{z}}{z-\overline{z'}}\right)^{k}
(\cosh^2r/2-\cosh^2\rho/2)^{-1/2}\times $ 
\begin{align} F\left(|k|,-|k|, 1/2, 1-\frac{\cosh^2r/2}{\cosh^2\rho/2}\right),\end{align}
and $\rho(z, z')$  is the distance between two points $z$ and $z'$ given \eqref{dist1}.
\end{cor}
\begin{proof}
The resolvent kernel and the heat kernel are related by the Laplace and Laplace inverse transforms as
\begin{align} G_k(\mu, w, w')=\int_0^\infty e^{\mu^2 t}H_k(t, w, w')dt,
\end{align}
or
\begin{align} H_k(t, z, z')=\frac{-1}{2 i\pi}\int_{c-i\infty}^{c+i\infty} e^{\mu t}G_k(\sqrt{\mu}, z, z')d\mu.
\end{align}
By combining the formula \eqref{key} and the formula  given by Prudnikov et al.  \cite{PRUDNIKOV et al.} p.52 
\begin{align}\label{Laplace-Transform} L^{-1}e^{-a\sqrt{p}}(x)=\frac{a}{\sqrt{4\pi x^3}}e^{-\frac{a^2}{4 x}},\ \ \  {\cal R}e p >0, \ \ \  {\cal R}e a^2 >0 ,\end{align}
and using Fubini theorem we get 
the formulas  \eqref{heat kernel} and the proof of the Corollary is finished.
\end{proof}
\section{Resolvent kernel with the  Morse potential on $\R$}
The Schr\"odinger operator with the Morse potential is given for $X\in \R$ by \eqref{Morse}
or equivalently for by setting $y=e^X \in \R^+$ by
\begin{align}\Lambda^{\lambda, k}_{\ln y}=\left(y\frac{\partial}{\partial y}\right)^{2}-2k\lambda y -\lambda^2 y^2.\end{align}
\begin{prop}
i) The modified Schr\"odinger operator with uniform magnetic field $\widetilde{{\cal D}}_{k}$ on the hyperbolic half plane  and  Schr\"odinger operator with the Morse diatomic molecular potential $\Lambda^{\lambda, k}_{\ln y}$ on the real line $\R$  are connected via the formulas
\begin{align}{\cal F}_x\left[y^{-1/2}{\cal D}^z_{k}y^{1/2}\Phi\right](\lambda, y)=\Lambda^{\lambda, k}_{\ln y}\left({\cal F}\Phi\right)(\lambda, y),\end{align}
where the Fourier transform is given by
\begin{align}
[{\cal F}f](\xi)=\frac{1}{\sqrt{2\pi}}\int_{\R}e^{-i x \xi}f(x)dx.
\end{align}
ii) The resolvent kernels  $\widetilde{G}_k(\mu, y, y')$ with uniform magnetic potential on the hyperbolic half plane and  $\widetilde{{\cal G}}_{\lambda,  k}(t, y, y')$  with the Morse potential are related by the formula
\begin{align}\label{connection}{\cal G}_{\lambda,  k}(\mu, y, y')= \frac{1}{\sqrt{y y'}}\int_{-\infty}^{\infty}e^{-i\lambda(x-x')}\widetilde{G}_k(\mu, z, z')d(x-x').\end{align}
\end{prop}
\begin{proof} The part i) is simple and in consequence is left to the reader. The part  ii) is a consequence of i).\\
\end{proof}
\begin{cor} The Schwartz kernel of the resolvent operator for the Schr\"odinger operator with Morse potential  
 is given by
\begin{align}\label{g} {\cal G}^{M}_{\lambda, k}(y, y', \mu)=\int^{\infty}_{0}e^{-i\mu b}W_{\lambda, k}(b, y, y') db,
\end{align}
where
\begin{align}\label{connection}W_{\lambda, k}(b, y, y') = \frac{1}{\sqrt{y y'}}\int_{-\infty}^{\infty}e^{-i\lambda(x-x')}\widetilde{W}_k(b, z, z')d(x-x').\end{align}
\end{cor}
\begin{proof}
This corollary is a consequence of the formulas \eqref{key} and \eqref{connection}.
\end{proof}

 Now, we shall prove the following theorem.
\begin{thm} The Schwartz kernel of the resolvent operator for the Schr\"odinger operator with Morse potential  
 is given by
\begin{align}\label{key-key} {\cal G}^{M}_{\lambda, k}(\mu, X, X')=\int^{\infty}_{0}e^{-\i\mu b}W_{\lambda, k}(b, y, y') db,
\end{align}
where
\begin{align}\label{w}W_{\lambda, k}(b, y, y')=C_{k}(4 y y')^{-|k|}\left(\frac{\partial}{\sinh (b /2) \partial b}\right)^{2|k|}
\nonumber\\ \frac{(2Z)^{4|k|}}{(Z+Y)^{2|k|}} e^{-i\lambda Z}\Phi_1(2|k|+1/2, 2|k|, 4|k|+1, 2i\lambda Z, \frac{2 Z}{Z+Y}),\end{align}
 $C_k=\frac{(-1)^{k}\Gamma(2|k|+1/2)}{2\Gamma(4|k|+1)\sqrt{\pi}}$,
 $Z=\sqrt{4y y'\cosh^2 (b/2)-(y+y')^2},$ \\ 
$Y=i\,sign(k)(y+y')$,
with the function $\Phi_1(a, b, c,  x, y)$ is the confluent hypergeometric function of two variables defined by the double series:(see for example 
Erdelyi et al.\cite{ERDELYI et al.} p. $225$).\\
\begin{align}\label{phi1}\Phi_1(a, b, c,  x, y)=\sum_{m, n \geq 0}\frac{(a)_{m+n}(b)_n}{c)_{m+n}m!n!}x^my^n,\end{align}
 for $|y|<1$ and its analytic continuation elsewhere.
\end{thm}
\begin{proof} Using the formula
$$(a^2+y^2)^{-1}=\int_0^\infty e^{- a x}\,\frac{\sin x y} {y}\, dx,$$
with $y=\sqrt{-\Lambda^{\lambda, k}}$ and $a=-i\lambda$, we can write
\begin{equation}\label{transmutation} (-\lambda^2-\Lambda^{\lambda, k})^{-1}=\int_0^\infty e^{- i\mu t}\,\frac{\sin t \sqrt{-\Lambda^{\lambda, k}}} {\sqrt{-\Lambda^{\lambda, k}}}\,dt,\end{equation}
comparing with the formulas \eqref{transmutation} and \eqref{key-key}, we have the result of the Corollary.
\end{proof}
\section{The time dependent Schr\"odinger equation under the Morse Potential on the real line $\R$} 
\begin{thm} For $k$ integer or half integer, the Cauchy problem  for the time dependent Schr\"odinger equation with
 the Morse Potential has the unique solution given by:
\begin{align} U(t, X)=\int_{-\infty}^\infty q_{\lambda, k}^M(t, X, X')f(X')dX',
\end{align}
where
\begin{align}\label{Heat-Kernel} q^M_{\lambda, k}(t, X, X')=\int_{|X-X'|}^<\infty> \frac{e^{-\frac{b^2}{4t} }}{(4\pi t)^{3/2}}W_{\lambda, k}^M (b, X, X')b db,\end{align}

$W_{\lambda, k}^M(b, X, X')=c_1(k)\left(\frac{d}{\sinh b/2 db}\right)^{2|k|}\\ \frac{Y^{4|k|}}{Z^{2|k|}}e^{(-k)(X+X')}e^{i\lambda z}\Phi_1(\alpha, \beta, \gamma, -2i \lambda Y, Z)$,\\
$c_1(k)=(-1)^{|k|+1}\frac{\Gamma(2|k|+1/2)}{2^{|k|}\Gamma(4|k|+1)\sqrt{\pi}},$\\
 $Y=2e^{(X+X')/2}\sqrt{\cosh^2 (b/2)-\cosh^2((X-X')/2)}$ and \\  $Z=\frac{2\sqrt{\cosh^2 (b/2)-\cosh^2((X-X')/2)}}{\sqrt{\cosh^2 (b/2)-\cosh^2((X-X')/2)}+i sign(k)\cosh((X-X')/2)},$\\
 $sign(k)=1$ if $k>0$ and $ sign(k)=-1$ if $k<0$.\end{thm}
\begin{proof}
The resolvent kernel and the heat kernel are related by the Laplace and Laplace inverse transforms as
\begin{align} {\cal G}^{M}_{\lambda, k}(X, X', \mu)=\int_0^\infty e^{\mu^2 t}q^{M}_{\lambda, k}(t, X, X') dt,
\end{align}
where
\begin{align} q^{M}_{\lambda, k}(t, X, X') =\frac{-1}{2 i\pi}\int_{c-i\infty}^{c+i\infty} e^{\mu t} {\cal G}^{M}_{\lambda, k}(X, X', \sqrt{\mu})d\mu.
\end{align}
By using the formulas \eqref{g}, \eqref{Laplace-Transform}, and the
 Fubuni theorem, we arrive at the formula \eqref{Heat-Kernel}.  The proof of the Corollary is finished.
\end{proof}

\section{Applications}
In this section, we introduce some applications.\\
{\bf  1- An integral representation of the product of two Whittaker functions.}\\
For $\alpha > 0$, $X> X' $, $\lambda k> 0$ the following formula holds
\begin{align}\label{Product}\frac{\Gamma(\alpha-k+1/2)}{\lambda\Gamma(1+2\alpha)}e^{-(x+y)/2}W_{k, \alpha}(2\lambda e^y)M_{k, \alpha}(2\lambda e^x)=\nonumber\\
\int^{\infty}_{0}e^{-\alpha b}W_{\lambda,k}^M(b, X, X')db\end{align}
where $W_{\lambda, k}$ is as in \eqref{w}.\\

Recall the the Schwartz kernel of resolvent of the Schr\"odinger operator with the Morse potential given by(Ikeda-Matsumoto \cite{IKEDA-MATSUMOTO} p. $83$)
\begin{align}\label{Resolvent-Morse}  G^{M}_{\lambda, k}(\mu, X, X')=\frac{\Gamma(\mu-|k|+1/2)}{\lambda\Gamma(1+2\mu)}e^{-(X+X')/2}W_{|k|, \mu}(2\lambda e^{X'})M_{|k|, \mu}(2\lambda e^X),\end{align}
where  $W_{k, \mu}(x)$ and $ M_{k, \mu}(x)$ are the Whittaker functions of the first and the second kind respectively.\\
Combining the formulas \eqref{Resolvent-Morse} and \eqref{key-key}, we obtain the formula \eqref{Product}.\\
{\bf 2 - An Integral formula} \\
Set\\
$ J(t, y, y')=\frac{\lambda\sqrt{2y y'}}{i(\pi^3 t)^{1/2}}\int^{\infty}_{0}\int^{\infty}_{0}\frac{\sinh \xi }{\sinh^2 u}\times $\\  $\exp{\left(-\lambda(y+y')\coth u -\frac{2\lambda\sqrt{y y'}\cosh\xi}{\sinh u}+2k u+2\frac{(\pi+i\xi)^2}{t}\right)} d\xi du$\\
For $\lambda > 0$, $k\in R$ and  $t>0$ , we  have \begin{align} \label{Integral-Formula}{\cal I} m [J(t, y, y')]=\int_{|X-X'|}^\infty \frac{e^{-\frac{b^2}{4t}}}{(4\pi t)^{3/2}}W_{\lambda, k}^M (b, y, y')b db.\end{align}
 where $W_{\lambda, k}$ is as in \eqref{w}.


To see the above formula  we recall the following formula for $q^M_{\lambda, k}(t, x, y)$ given in Alili et al.\cite{ALILI et al.}
\begin{align}\label{Morse-Heat-Alili} q^M_{\lambda, k}(t, x, y)=\int^{\infty}_{0}e^{2ku}\frac{1}{2\sinh u}\exp{\left(-\lambda(e^x+e^y)\coth u\right)}\theta_{\bar{\Phi}}(t/4)du,\end{align}
where $$\theta_r(t)=\frac{r}{(2\pi^3 t)^{1/2}}e^{\pi^2/2t}\int^{\infty}_{0}e^{-\xi^2/2t}e^{-r\cosh \xi}\sinh \xi \sin(\pi\xi/t)d\xi,$$
and
$$\bar{\Phi}=\frac{2\lambda\exp{(x+y)/2}}{\sinh u}.$$
In view of the uniqueness of the solution of the time dependent  Schr\"odinger equation with the Morse potential we obtain the formula\eqref{Integral-Formula} by comparing \eqref{Morse-Heat-Alili} and \eqref{Heat-Kernel}.
 Note that using formulas Magnus et al. \cite{MAGNUS et al.} p.$305$,
$$W_{0, \alpha}(z)=\sqrt{z/\pi}K_\alpha(z/2),\ \ \ 
M_{0, \alpha}(z)=2^{2\alpha}\Gamma(\alpha+1)\sqrt{z}I_\alpha(z/2),$$
where $I_\nu$ and $K_\nu$ are the Modified Bessel functions of the first and the second kind
and   (\cite{MAGNUS et al.} p. 283)
$$\Phi_1(\alpha, 0, \gamma, x, y)={}_1F_1(\alpha, \gamma, x), \ \  \
F(\nu+1/2, 2\nu+1, 2 i z)=\Gamma(1+\nu)e^{i z}(z/2)^{-\nu}J_{\nu}(z),$$
with ${}_1F_1$ and $J_{\nu}$ are respectively the confluent hypergeometric and the Bessel functions of the first kind, 
 we have for $k=0$, 
\begin{align}\label{0} {\cal G}^{M}_{\lambda,  0}(\mu, X, X')=\int^{\infty}_{0}e^{-i\mu b}W_{\lambda, 0}(b, y, y') db,
\end{align}
$$W_{\lambda, 0}(b, y, y')=\frac{1}{2}J_0(|\lambda|\sqrt{2 y y' \cosh b-y^2-y'^2}).$$ 
The formula \eqref{Product} reduces to the following integral representation of the product of the Bessel
functions
Lebdev\cite{LEBEDEV} p. 140
$$I_\mu(u)K_\mu(v)=(1/2)\int^{\infty}_{0}e^{-i\mu b}J_0(\sqrt{2 u v\cosh b-u^2-v^2})db, u >0, v>0.$$

  M.V. Ould Moustapha, \textsc{Department of Mathematics,
 College of Arts and Sciences-Gurayat,
 Jouf University-Kingdom of Saudi Arabia }.\\
\textsc{Facult\'e des Sciences et Techniques
Universit\'e de  Nouakchott Al-asriya,
Nouakchott-Mauritanie.}\\
  \textit{E-mail address}: \texttt{mohamedvall.ouldmoustapha230@gmail.com}
  \end{document}